OpenStudy (anonymous):

integral of (t^2 + 64) / (t^2 - 64)?

OpenStudy (perl):

i would do long division first

OpenStudy (anonymous):

\[\int\limits\limits_{}^{} \frac{ (t^2 + 64) }{ (t^2 - 64) }\]

OpenStudy (anonymous):

i did long division then I got \[1 + \frac{ 128 }{ t^2 - 64 }\]

OpenStudy (perl):